OpenStudy (anonymous):

A ball is thrown up into the air such that its height (h) from the starting point is given by h(t) = 50t – 5t^2, where t is the time (in seconds) elapsed from the moment the ball is released. During which time interval will the height of the ball be 80 meters or higher above the starting point?

OpenStudy (anonymous):

@andrewhaze @jackmullen55

OpenStudy (anonymous):

@nessachole02 @PizzaLover123

OpenStudy (anonymous):

@taylorswift131313 @ZnappyDooZ

OpenStudy (anonymous):

So what you want to do is input different values for t to find what makes the equation equal to 80. I'll do an example

OpenStudy (anonymous):

Lets say that the variable t is equal to 2

OpenStudy (anonymous):

\[80(2) = 50(2) – 5(2)^2\]

OpenStudy (anonymous):

so if I solve this equation ill get the answer?

OpenStudy (anonymous):

Simplify this and you get \[80 = 100 - 5(4) = 100 - 20 = 80\] Ok well I just accidentally solved it for you

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

The answer is t = 2

OpenStudy (anonymous):

that's not one of my choices though that's what i don't get

OpenStudy (anonymous):

What are the choice you didn't specify

OpenStudy (anonymous):

A) ( 3,6) B) ( 4,10) C) (2,8) D) ( 2,6)

OpenStudy (anonymous):

It's C

OpenStudy (anonymous):

thanks but where did u get the 8?

OpenStudy (anonymous):

I had the same question in a test once. It was (2, 8)

OpenStudy (anonymous):

never mind i got it

OpenStudy (anonymous):

Yeah